Fermi National Laboratory


Accelerator Update for January 17, 2005 - January 19, 2005

Monday January 17
The day shift began with the Tevatron (TeV) in shot setup, with MiniBooNE taking beam, but with the Meson experiment MCenter off due to shot setup.

Operations established store 3925 at 11:23 AM with an initial luminosity of 104.96E30. This store set a NEW RECORD LUMINOSITY. The Antiproton Source (Pbar) resumed stacking and MCenter resumed taking beam.

Operations put the Recycler into shot setup for antiproton transfers at 4:49 PM. By 6:10 PM, the Recycler had stashed six shots from Pbar; one transfer failed. Pbar resumed stacking.

Operations reported that around 7 PM Pbar set a NEW STACKING RECORD with a rate of 14.94mA per hour.

At 7:48 PM, a Main Injector (MI) RF station (MIRF05) tripped off. Operators bypassed the station since they couldn't reset it.

At 8:04 PM, a Pbar power supply (PMAG) tripped off. After experts were consulted, Pbar resumed stacking about twenty minutes later.

Operations began setting up beam for MCenter at 8:06 PM.

Operations put the Recycler into shot setup at 10:15 PM. By 11:02 PM, they had stashed five transfers of antiprotons. Pbar resumed stacking.

Operations reestablished beam to MCenter and established beam to MTest at 11:33 PM.

Tuesday January 18
The midnight shift began with Operations monitoring store 3925 and the Recycler stash, with Pbar stacking, and with MiniBooNE, MCenter, and MTest taking beam.

At 3:38 AM, MIRF18 tripped off for the first of four times over three and a half hours. It reset okay each time.

At 9:23 AM, Operations heard that a water main had broken by the TeV sector E3 Pond. Operators turned off the E3 pond pump and notified experts. As you can see from the pictures below, the water bubbled up from under the road and froze. Roads and Grounds workers salted the road.

At 11:31 AM, Operations received a report of a problem with the Neutron Therapy Facility's (NTF) 32-degree bend magnet power supply. A power supply technician investigated and found a bad cable termination. The work took about three and half hours to complete.

Operations turned off the Linac beam for an access at 4:14 PM. This access was to further investigate the NTF cable problem. This also shut off all downstream beam.

The experts came out of the Linac access at 5:12 PM and reported that cables in the tunnel also had to be reterminated. The Run Coordinator scheduled this work for Thursday morning.

Linac beam resumed at 5:15 PM, which meant that Pbar resumed stacking and MiniBooNE, MTest, and MCenter resumed taking beam.

Operations terminated store 3925 at 7:02 PM and put the TeV into shot setup at 7:56 PM.

NuMI completed testing its horn at 9:08 PM.

At 9:26 PM, Pbar had three BPMs stop working at Debuncher house 31. Operators notified the experts.

At 9:44 PM, Operations reported that the TeV aborted on the seventh transfer of protons. Operations reloaded the protons.

At 10:44 PM, a Booster quadrupole power supply (B:QDB3) tripped off for the first of four times within seven minutes. It would not ramp after the last trip.

Operations established store 3926 at 10:59 PM with an initial luminosity of 104.25E30. Pbar resumed stacking.

Operators had B:QDB3 back on by 11:18 PM.

Wednesday January 19
The midnight shift began with Operations monitoring store 3926, with Pbar stacking, and with MiniBooNE taking beam.

At 12:08 AM, a Linac RF station (LRF1) went to zero with no trip status. Operators managed to bring the station up locally.

Earlier, the MCenter interlocks had dropped. At 1:22 AM, operators search and secured the enclosure and then began preparing the experiment for beam.

At 1:25 AM, MIRF11 tripped off and wouldn't reset. Operators bypassed the station.

At 1:54 AM, Operations reported that they were still having problems with B:QDB3.

At 2:14 AM, problems with the I- Source were causing problems with MiniBooNE. Operators tried to fix the problem with no success. They called in an expert.

Operations put the Recycler into shot setup at 4:32 AM.

The Linac expert reported at 6:21 AM that the I- Source problem was due to high conductivity. He made a small temporary adjustment.

The Recycler stash was completed and stacking resumed at 7:17 AM.

MCenter resumed taking beam at 7:55 AM.

The Plan for Wednesday
The plan for today is to keep the store in the machine throughout the night and during the access on Thursday. This will be a six-hour access into Linac and Booster beginning at 7 AM.


Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.



last modified 1/19/2005   email Fermilab